Date: Thu, 22 Oct 1998 18:08:41 -0700 (PDT)
To: 2mass
Subject: IPAC 2MASS WG Meeting #169 Minutes
Cc: chas, stiening, bgreen
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: h5JR0cG3DLuyJIYvY7Gf2A==

           IPAC 2MASS Working Group Meeting #169 Minutes

Attendees: R. Cutri, S. Van Dyk, R. Beck, T. Evans, R. Hurt, 
           H. McCallon, S. Wheelock, R. Tam, W. Wheaton, 
           T. Jarrett, T. Handley, D. Kirkpatrick


1.) Project Update 

2.) New and Standing Action Items 


1.) Project Update 

     R. Cutri spent a night at Mt. Hopkins recently.  He reported that the night
consisted of a frenzy at dusk, a lull for the rest of the night (as the 
telescope automatically observes), and a renewed frenzy at dawn.

     It appears that a problem occurred with the previous night's data (981020n)
in the north.  A strange pattern appears in the images, indicating some problem
with the camera electronics.  This marks the end of the use of the old Gatir 
electronics.  R. Stiening was headed to Mt. Hopkins to swap in the new Leach
electronics.  Data from the facility *are* relatively indistinguishable between
old and new electronics, but, obviously, the data taken with the new electronics
will have to be characterized.  Some work will therefore be involved at IPAC.

     Everyone has submitted their `wish lists' for 2MAPPS 3.0.  Also, WG members
have submitted their performance reports, which will become a monthly

     The `oddity of the week' appears to be hot pixels in the southern data.
     B. Nelson will be moved to a regular staff position soon and will serve as
an additional QA scientist.  That means that QA should become an additional 33%
more efficient.  However, this is balanced (precariously) against the fact that
in a week, dino (aka SOB) will be going online, leading to 50% more processing
efficiency.  Dino will be loaded with 200 Gb of disk.  Dino likely represents
the last of the Enterprise 450's that will be purchased.

     Since the recent recollimation of the northern telescope, the standard
second image moment ratio is no longer a sensitive diagnostic for image
elongation in the J and H bands.  A more sensitive diagnostic, developed by T. 
Jarrett, evaluates the true aspect ratio at the appropriate position angle of 
the elongation of the image.  The new diagnostic will be reported in the 
standard qDATEh.psf.see file, and that file will have a slightly different 
format, with three extra columns for the position angles.  This new routine must
be tested further, but R. Cutri would like it implemented in the QA process by 
week's end.

     S. Wheelock would like to implement "PSFCOR", the routine to interpolate
the photometric cross-scan bias correction across cal scans.  R. Cutri had not
yet had a chance to read her memo on the routine.

     T. Evans reported that the database came back up after the weekend's power
outage.  She is still validating the database load that took place over the
weekend.  However, it is clear that one of the southern database indices got
corrupted.  This could just be the result of a transient in the system.  The
server may have gotten into a weird state, so caveat emptor to all team members.
The Sampler seems to be OK, but this had just gotten reloaded.  She wants to
reload the point source catalog for the Sampler.

     T. Evans and T. Handley reported that a recent reindexing of the database
locked up OPS for 25 hours, which is too long.  Only one-tenth of the sources
observed are currently in the database (only two nights of southern data are
currently in the database), which raises the horrific concern whether future 
database loads and indexing will take geometrically longer and forbidding 
amounts of time.  Indexing is clearly needed for good queries to be 
accomplished, however, not if the indexing takes weeks to perform! 

     H. McCallon produced a tome on the evaluating and updating of positional
uncertainties.  These documents have now been placed online at the 2MASS 
internal website.  The issues remain whether old positions plus deltas are
put into the source catalogs, or new positions are added to the catalogs, once
the old positions are corrected.  The positional uncertainties need to be 
updated for the spring data release.  Team members are encouraged to peruse
his memos online.

     W. Wheaton asked about information that can be obtained about when the
cameras warm up and how this affects the detector saturation levels in the 
north.  R. Cutri urged W. Wheaton to first get the saturation levels for the 
south and then possibly reanalyze the saturation levels for the north later.

     T. Jarrett asked about "J-banding", where J jumps are seen along half of
the coadds in the south.  He pointed out that these jumps could kill galaxy
photometry in GALWORKS.  The jumps could lead to 20%, or more, errors in the
photometry.  R. Tam will look at the amplitudes of the jumps and how many scans
they affect.  These could be bias jumps, similar to the "J quadrant dropouts"
seen in the north.  So, a precedent exists for this problem.  The solution may
be to have to throw out these data.  Individual coadds must therefore be
examined first.  The problem has only been detected so far in the early southern
nights that have been processed.  R. Beck will process some sci scans in recent
nights, and nights in between, so that it can be determined whether or not the
problem has persisted since March/April.

2.) New and Standing Action Items

%                                                      %
%                 ACTION ITEMS                         %
%                                                      %

      R. Cutri did not assign any new action items.  [However, in an e-mail
since the WG meeting, he has set a deadline of Dec 2 for the performance
reports covering October and November.]

(Standing Action Items:)

a) Team members should continue testing the IRSA tools.

b) Team members should still contemplate the sci-staff structure proposals and 
voice their opinions.

c) Team members should continue to analyze data for spring data release.